OpenStudy (moses238):

The sum of two consecutive integers is at least 46. What is the least possible pair of integers?

OpenStudy (anonymous):

divide 46 by 2, whatever you get is the smaller number

OpenStudy (anonymous):

because it has to be at least 46

OpenStudy (moses238):

thanks

OpenStudy (anonymous):

ur welcome :)

OpenStudy (moses238):

hold on

OpenStudy (moses238):

the answers they gave me were: a. 21 and 23 b. 23 and 24 c. 22 and 23 d. 24 and 25

OpenStudy (anonymous):

Its B. Pretty self evident, right?

OpenStudy (moses238):

why

OpenStudy (ybarrap):

Let the first integer be \(a\) and the second be \(b=a+1\) then $$ a+b=a+a+1\ge 46\\ 2a+1\ge 46\\ 2a\ge45\\ a\ge \cfrac{45}{2}=22.5\\ \implies a=23,b=24 $$ Make sense?